# Taking the Garbage Out of Data-Driven Prediction Across Climate Timescales

### v1

@@ 1 Introduction
The integration of artificial intelligence and machine learning (AI/ML) in weather and climate science is rapidly revolutionizing predictions and our understanding of Earth climate system. Such techniques offer increased capabilities in handling large datasets, identifying complex patterns, and making accurate predictions. Recent attention has been heavily centered on model choice (Dueben and Bauer, 2018; de Burgh-Day and Leeuwenburg, 2023; Molina et al., 2023b) and explainability of the predictions (e.g. Mamalakis et al., 2022, 2023; Yik et al., 2023; Camps-Valls et al., 2025). However, the effectiveness of data-driven models strongly depends on data quality. This consideration is paramount and led to the popular adage “garbage in, garbage out,” credited to computer programmers in the late 1950s (Lidwell et al., 2003). Simply put, if flawed or poor-quality data are fed into a model, the resulting predictions will likely also be flawed.
@@
Data quality can be evaluated in multiple ways. One method is based on sample size and the fidelity of the data. These considerations are important for AI/ML applications and have been addressed in several works across disciplines (e.g., Dueben et al., 2022; de Burgh-Day and Leeuwenburg, 2023; Zantvoort et al., 2024; Xie et al., 2025). For the atmospheric sciences, so-called “benchmark platforms” that provide datasets ready for AI/ML applications have been developed – e.g., WeatherBench (Rasp et al., 2020, 2024), AQ-Bench (Betancourt et al., 2021), ClimSim (Yu et al., 2024), and ChaosBench (Nathaniel et al., 2024). However, even if one has high quality data with sufficient samples, a data-driven model may still produce poor results, as “garbage” can result from other erroneous assumptions or treatments of the input data: e.g., wrong assumptions around data biases, incorrect assessment of the “true” distribution, incorrect classification labels, and inconsistent thresholds and definitions of phenomena.
@@
Climate data present unique challenges for use in data-driven models. Most climate datasets are inherently spatiotemporal, sparse, and possess spatial and temporal autocorrelations. The data are often nonstationary, especially in recent decades, due to anthropogenic climate change. This effect changes the core statistics of the variables of interest (e.g., temperature, wind, cloud cover, geopotential height). Studies handle nonstationarity in different ways. For example, when removing trends from a time series, some studies may simply remove a linear or second-order polynomial trend, while others may use more complex techniques, such as empirical mode decomposition (e.g., Huang et al., 1998), to detect changing trends over time. Along with changing background statistics, climate variables also have varying distributions, many of which are non-normal (e.g., gamma, bimodal, log-normal, and skew-normal), and exhibit non-linear interactions among themselves. As such, many traditional methods in statistics and AI/ML cannot simply be used “out of the box” when working with climate variables. Additionally, climate data are collected from diverse sources, including satellite observations, weather stations, and climate models, and can be noisy, incomplete, and heterogeneous. The rapid development of AI/ML applications in climate prediction necessitates a guiding set of recommended preprocessing steps for climate data to secure some degree of harmony between studies and applications. As such, understanding the rationale for why certain data preprocessing steps are taken can improve trust in these methods for end users by demystifying the process and providing ways to evaluate and critique the models and their results.
@@
The aim of this paper is to present recommended practices on proper data preprocessing steps aimed for climate (specifically, subseasonal to decadal and longer) prediction studies using AI/ML. This paper has three overarching goals:
@@ 2 Recommendations for Initial Steps in Data Preparation and Minimizing Data Leakage
The initial step in climate prediction is to clearly identify what the researcher wants to predict, over which timescale(s) to make this prediction, and what AI/ML methods are most appropriate for addressing the prediction problem. This step is essential and should be carefully considered –- the reader is referred to other works for discussion of problem setup and different AI/ML methods used in prediction studies (e.g., Molina et al., 2023b; Yang et al., 2024; Camps-Valls et al., 2025). Upon making these decisions, the next step is to select a set of potential input features (i.e., predictors) and outputs used to generate a prediction or classification with that AI/ML model during training or testing. Most prediction problems will use a supervised learning framework requiring inputs and outputs; unsupervised learning applications only need inputs. Inputs and outputs can take the form of a numerical value (e.g., the Niño 3.4 index), numerical data fields (e.g., sea surface temperature), categories (e.g., an El Niño or La Niña event), or a probability distribution (e.g., ensemble spread). More about the terminology above can be found in Chase et al. (2022).
@@
We recommend a period of “data exploration” first to identify key first-order statistics of the input features and locate any missing or erroneous data. As mentioned, climate variables can possess autocorrelation and covariances between them in space and time. Quantifying these relationships allows us to identify the effective sample size $\left(N_{\mathrm{eff}}\right)$, which can often be smaller than the total number of samples ($N$; e.g., Bretherton et al., 1999). Recognizing this concept is important, as it will impact the choice for the number of training samples needed. Sparse data (e.g., weather stations) introduce additional challenges, such as potential sampling biases. Interpolation can be used to fill undersampled data, but care must be taken with the choice of interpolation method (e.g., bilinear vs. piecewise constant). Finally, understanding the distributions and trends of the variables will also inform the preprocessing steps needed (see Section 3).
@@
After initial data exploration, one should establish a representative training dataset, which will be used to “fit” the model’s parameters (e.g., weights or coefficients). In choosing the number of training samples needed, one should consider: (a) model complexity, (b) available computational resources, (c) number of total input features and their ability to represent a range of possible outcomes, and (d) required/or desired accuracy. Having many input features, however, requires a larger training sample size for the given model. As such, input features can be reduced using filtering methods (i.e., eliminating features with little-to-no statistical relationship with the target) or feature extraction (i.e., dimensionality reduction using, for example, principal component analysis). We advise considering one or more of these methods for feature selection for one’s AI/ML problem.
@@
Thereafter, data “splits” need to be decided. For this, the datasets are typically divided into three distinct subsets: (1) training (used for fitting the model); (2) validation (used for adjusting configurable components of the model, known as hyperparameters), and (3) testing (used for final model evaluation). Commonly used ratios for training, validation, and testing sets include 60:20:20 or 80:10:10. Since the validation dataset is used during model development, it should not be considered in assessing test results. Doing so contributes to data leakage (discussed below), potential data misuse, and even unethical use of AI/ML. We further recommend that results for the training, validation, and testing datasets be openly reported to assess, for example, the generalizability properties of the model (e.g., how well the model will perform with the same predictors at a future time). We encourage the use of cross-validation (CV) methods to mitigate learned bias as well as using the full training and validation sets for more robust hyperparameter tuning (e.g., Sweet et al., 2023). $k$-fold cross-validation is the most commonly used cross-validation method in practice; Table 1 shows other CV methods, along with their limitations.
@@
Leave-one-out Suitable for small data. Each sample serves as validation once. Computationally expensive for large data.
@@
As mentioned, improper splitting choices of the data can inadvertently lead to data leakage, resulting in artificially inflated performance metrics and poor generalization of the results. Leakage of training or testing data occurs when input features contain or are derived from information that can reveal the target variable and would not be (feasibly) available during real-time predictions (e.g., predicting tomorrow’s total precipitation with today’s total precipitation). Data leakage is especially problematic in climate prediction, where variables often possess strong temporal autocorrelation (i.e., “memory”), quasi-periodicity, and/or long-term trends. Figure 1a illustrates an example of data leakage in detrending a time series of monthly-mean temperatures. The use of the full-time period for detrending (1940–2024; Fig. 1a, red line) versus just using the trend during the training period (1940–2010; Fig. 1a, black line) results in much cooler temperatures during the test period (2011–2024), biasing the AI/ML model results. Test data leakage into the training set can also occur during the scaling and normalizing of the input features. Considering temporal autocorrelation in climate data, we recommend block splitting; i.e., the dataset is split into temporally continuous training/validation and testing blocks or periods (e.g., Fig. 1a), with a gap between the two blocks adequately long relative to the timescales of interest to increase independence (e.g., Zhu et al., 2023).
@@
Figure 1: Illustration of various time series transformations using ERA5 (Hersbach et al., 2020). Data leakage is shown in panel (a), where the black line represents no data leakage due to detrending using only the training set, and the pink line represents training and testing data leakage due to detrending using the full data period. Data leakage results in 54% cooler temperatures during the testing period. Various kernel density estimations of data transformations are shown in panel (b) for a non-normal dataset (black dashed line).
@@
To prevent leakage, we recommend splitting data before any preprocessing steps. The testing data should be set aside and not used until model development is finalized. For instance, feature selection and calculation of scaling factors or trends should be made with the training/validation data after the split and later applied to the testing data if necessary. For data with strong autocorrelation, quasi-periodicity, and/or trends, data splitting by obeying time series dependencies or phenomena properties (e.g., choosing El Niño “years” based on season instead of calendar year) will prevent artificially inflated skill metrics. Since subseasonal and longer temporal dependencies can result in an insufficient $N_{\mathrm{eff}}$ to represent the processes of interest (e.g., Mayer et al., 2024), other techniques should be considered. For example, using data from long simulations of climate models to develop the initial versions of the AI/ML models and overcome the limitations of short observational records is becoming more common (e.g., Ham et al., 2019; Rivera Tello et al., 2023). This practice assumes that climate models adequately represent the phenomena of interest and that fine-tuning procedures can correct existing biases. To address spatial autocorrelation, one can stratify geographically distinct regions and build buffer zones between them. If regions are climatologically distinct and such properties are important for model performance, ensuring that data splits contain samples from respective regions while limiting temporal autocorrelation may be appropriate.
@@ 3.1 Working with numerical-valued data: Anomalies and standardization
Most problems in climate prediction involve predicting the anomaly (i.e., departure from an average) in a given field, allowing for skill evaluation beyond just the varying climatology (Hamill and Juras, 2006). However, computation of anomalies depends on the problem design, and the computation method can yield different interpretations. A baseline period may be fixed (e.g., 1991–-2020) or centered rolling windows (e.g., $\pm 15$ years), the latter of which may address nonstationarity. Choosing a baseline period early in the record may result in artificial skill gains during evaluation due to the effects of climate change (Wulff et al., 2022). Choosing a base period later in the record, particularly one with a strong trend, will skew early values. Long-term trends should also be removed to avoid overinflating correlations between input features or relationships derived therein. When calculating anomalies and trends, the spatial element of the climate variables should also be considered. These quantities can be computed using global, regional, or grid cell-specific climatology. A global or regional baseline climatology (latitudinally weighted if using gridded data) may be necessary when spatial anomaly patterns must be preserved (e.g., modes of variability). A grid cell-specific climatology accounts for local variations, which may be useful when interested in strong horizontal gradients (e.g., precipitation anomalies). Altogether, we recommend that there should not be a “one-size-fits-all” approach to anomaly calculation.
@@
A special case exists when using hindcast simulations (i.e., forecasts made with a model for past events). Relatively short hindcast databases (e.g., 10-20 years long) restrict choices for climatologies, potentially leading to averages that use “future” periods and artificially inflating prediction skill (Risbey et al., 2021). Furthermore, hindcast climatologies are special because they are a function of the initialization date and lead time, so as to avoid model drift and bias. Therefore, when working with hindcast simulations, we suggest computing separate lead time-based climatologies and applying smoothing windows to multiyear averages to address gaps in initialization dates (e.g., Pegion et al., 2019). For ensemble-based hindcasts, the above-suggested recommendations for anomaly and trend calculations work well for single-model analyses. When considering multi-model means, anomalies should be computed for each model separately.
@@
Finally, feature scaling, either normalization (e.g., min-max scaling) or standardization (i.e., $z$-scoring), of data inputs for AI/ML applications must be carefully done. Single fields with large magnitudes and/or multiple fields with varying magnitudes can cause instability and prevent convergence, requiring feature scaline. We recommend that users consistently scale all features while also considering model design when making such a choice. Feature scaling is unnecessary when using tree-based models (e.g., random forests) since they use feature thresholds in their structure but is especially important when using algorithms sensitive to variance, data containing outliers, and when using distance-based models (e.g., $k$-means clustering). Since outliers can disproportionately affect feature scaline, unless one is interested in predicting outlier events, we suggest either removing the outliers or winsorizing the distribution – i.e., outlier values are reassigned to a specified percentile of the data.
@@
When dealing with variables that have non-Gaussian distributions, other preprocessing transformations can help align input features and should be performed before feature scaling. Figure 1b provides examples of different transformations done on the distribution of monthly-mean total precipitation maxima in the Indian Monsoon region, depending on user needs: (a) the log transformation (Fig. 1b, red line), (b) the Box-Cox (Fig. 1b, solid blue line; Box and Cox, 1964) or Yeo-Johnson (Fig. 1b, dashed red line; Yeo and Johnson, 2000) transformations, and (c) quantile transformations (Fig. 1b, light blue and solid red lines). Feature scaling should then be applied after these transformations.

### v2

# Setting the Standard: Recommended Practices for Data Preprocessing in Data-Driven Climate Prediction

@@ 1 Introduction
The integration of artificial intelligence and machine learning (AI/ML) in weather and climate science is rapidly revolutionizing predictions and our understanding of Earth climate system. Such techniques offer increased capabilities in handling large datasets, identifying complex patterns, and making accurate predictions. Recent attention has been heavily centered on model choice (i.e., selecting which type of ML model is most appropriate; dueben2018; deburgh-day2023; molina2023review) and explainability of the predictions (e.g. mamalakis2022; mamalakis2023; yik2023; camps-valls2025). However, the effectiveness of data-driven models strongly depends on data quality. This consideration is paramount and led to the popular adage “garbage in, garbage out,” credited to computer programmers in the late 1950s (lidwell2003). Simply put, if flawed or poor-quality data are fed into a model, the resulting predictions will likely also be flawed.
@@
Data quality can be evaluated in multiple ways. One method is based on sample size and the fidelity of the data. These considerations are important for AI/ML applications and have been addressed in several works across disciplines (e.g., dueben2022; deburgh-day2023; zantvoort2024; xie2025). For the atmospheric sciences, so-called “benchmark platforms” that provide datasets ready for AI/ML applications have been developed – e.g., WeatherBench (rasp2020; rasp2024), AQ-Bench (betancourt2021), ClimateBench (watson-parris2022), ClimSim (yu2024), and ChaosBench (nathaniel2024). However, even if one has high quality data with sufficient samples, a data-driven model may still produce poor results, as “garbage” can result from other erroneous assumptions or treatments of the input data: e.g., wrong assumptions around data biases, incorrect assessment of the “true” distribution, incorrect classification labels, and inconsistent thresholds and definitions of phenomena.
@@
Climate data present unique challenges for use in data-driven models. Most climate datasets are inherently spatiotemporal, sparse, and possess spatial and temporal autocorrelations. The data are often nonstationary, especially in recent decades, due to anthropogenic climate change. This effect changes the core statistics of the variables of interest (e.g., temperature, wind, cloud cover, geopotential height, sea level rise). Studies handle nonstationarity in different ways. For example, when removing trends from a time series, some studies may simply remove a linear or second-order polynomial trend (e.g. Long2025), while others may use more complex techniques, such as empirical mode decomposition (e.g., huang1998), to detect changing trends over time. Along with changing background statistics, climate variables also have varying distributions, many of which are non-normal (e.g., gamma, bimodal, log-normal, and skew-normal), and exhibit non-linear interactions among themselves. As such, many traditional methods in statistics and AI/ML cannot simply be used “out of the box” when working with climate variables. Additionally, climate data are collected from diverse sources, including satellite observations, weather stations, and climate models, and can be noisy, incomplete, and heterogeneous. The rapid development of AI/ML applications in climate prediction – i.e., forecasting the state of the climate system (in probabilities) on timescales ranging from several weeks to a couple of decades in the future – necessitates a guiding set of recommended preprocessing steps for climate data to secure some degree of harmony between studies and applications. As such, understanding the rationale for why certain data preprocessing steps are taken can improve trust in these methods for end users by demystifying the process and providing ways to evaluate and critique the models and their results.
@@
The aim of this paper is to present recommended practices on proper data preprocessing steps aimed for climate prediction studies using AI/ML. This paper has three overarching goals:
@@ 2 Recommendations for Initial Steps in Data Preparation and Minimizing Data Leakage
The initial step in climate prediction is to clearly identify what the researcher wants to predict, over which timescale(s) to make this prediction, and what AI/ML methods are most appropriate for addressing the prediction problem. This step is essential and should be carefully considered –- the reader is referred to other works for discussion of problem setup and different AI/ML methods used in prediction studies (e.g., molina2023review; yang2024; camps-valls2025). Upon making these decisions, the next step is to select a set of potential input features (i.e., predictors) and outputs used to generate a prediction or classification with that AI/ML model during training or testing. Most prediction problems will use a supervised learning framework requiring inputs and outputs; unsupervised learning applications only need inputs. Inputs and outputs can take the form of a numerical value (e.g., the Niño 3.4 index), numerical data fields (e.g., sea surface temperature), categories (e.g., an El Niño or La Niña event), or a probability. More about the terminology above can be found in chase2022.
@@
We recommend a period of “data exploration” first to identify key first-order statistics of the input features and locate any missing or erroneous data. As mentioned, climate variables can possess autocorrelation and covariances between them in space and time. Quantifying these relationships allows us to identify the effective sample size $\left(N_{\mathrm{eff}}\right)$, which can often be smaller than the total number of samples ($N$; e.g., bretherton1999). Recognizing this concept is important, as it will impact the choice for the number of training samples needed. Sparse data (e.g., weather stations) introduce additional challenges, such as potential sampling biases. Interpolation or imputation can be used to fill undersampled data, but care must be taken with the choice of method (e.g., bilinear vs. piecewise constant). Finally, understanding the distributions and trends of the variables will also inform the preprocessing steps needed (see Section 3).
@@
After initial data exploration, one should establish a representative training dataset, which will be used to “fit” the model’s parameters (e.g., weights or coefficients). In choosing the number of training samples needed, one should consider: (a) model complexity, (b) available computational resources, (c) number of total input features and their ability to represent a range of possible outcomes, and (d) required/or desired accuracy. Having many input features, however, often requires a larger training sample size for the given model. As such, input features can be reduced using filtering methods (i.e., eliminating features with little-to-no statistical relationship with the target) or feature extraction (i.e., dimensionality reduction using, for example, principal component analysis). We advise considering one or more of these methods for feature selection for one’s AI/ML problem.
@@
Thereafter, data “splits” need to be decided. For this, the datasets are typically divided into three distinct subsets: (1) training (used for fitting the model); (2) validation (used for adjusting configurable components of the model, known as hyperparameters), and (3) testing (used for final model evaluation). Commonly used ratios for training, validation, and testing sets include 60:20:20 or 80:10:10. Since the validation dataset is used during model development, it should not be considered in assessing test results. Doing so contributes to data leakage (discussed below), potential data misuse, and even unethical use of AI/ML. We further recommend that results for the training, validation, and testing datasets be openly reported to assess, for example, the generalizability properties of the model (e.g., how well the model will perform with the same predictors at a future time). We encourage the use of cross-validation (CV) methods to mitigate learned bias as well as using the full training and validation sets for more robust hyperparameter tuning (e.g., sweet2023). $k$-fold cross-validation is the most commonly used cross-validation method in practice; Table 1 shows other CV methods, along with their limitations.
As mentioned, improper splitting choices of the data can inadvertently lead to data leakage, resulting in artificially inflated performance metrics and poor generalization of the results. Leakage of training or testing data occurs when input features contain or are derived from information that can reveal the target variable and would not be (feasibly) available during real-time predictions (i.e., taking information from the future, which one would not have access to at the time of prediction). Data leakage is especially problematic in climate prediction, where variables often possess strong temporal autocorrelation (i.e., “memory”), quasi-periodicity, and/or long-term trends. Figure 1a
@@
Leave-one-out Suitable for small data. Each sample serves as validation once. Computationally expensive for large data. Does not address temporal dependencies.
@@
illustrates an example of data leakage during detrending of a global (latitude weighted) time series of the monthly average of daily maximum temperatures. The use of the full-time period for detrending (1940–2024; Fig. 1a, red line) versus the training period (1940–2010; Fig. 1a, black line) results in much cooler ground-truth (i.e., target) temperatures during the (hypothetical) test period (2011–2024), which can potentially bias the AI/ML model. The choice of a linear (1-degree) or quadratic (2-degree) polynomial for detrending can also have notable effects, where the 2nd-degree polynomial minimizes data leakage effects during the (hypothetical) test period and removes low-frequency artifacts in the training period (Figs. 1a and b). Data leakage can also occur
@@
Figure 1: Illustration of various time series preprocessing transformations using ERA5 (hersbach2020). Preprocessing data leakage for global (latitude weighted) monthly mean of daily maximum temperature anomalies derived from a monthly climatology (1981-2010) is shown in panels (a) and (b), where the black lines represent no data leakage due to detrending using only the training set period, and the pink lines represent leakage due to detrending using the training and (hypothetical) testing periods. Various kernel density estimations of data transformations are shown in panel (c) for a non-normal precipitation variable (black dashed line).
during the scaling and normalizing of the input features. Considering temporal autocorrelation in climate data, we recommend block splitting; i.e., the dataset is split into temporally continuous training/validation and testing blocks or periods (e.g., Fig. 1a), with a gap between the two blocks adequately long relative to the timescales of interest to increase independence (e.g., zhu2023).
@@
To prevent leakage, we recommend splitting data before any preprocessing steps. The testing data should be set aside and not used until model development is finalized. For instance, feature selection and calculation of scaling factors or trends should be made with the training/validation data after the split and later applied to the testing data. For data with strong autocorrelation, quasi-periodicity, and/or trends, data splitting by obeying time series dependencies or phenomena properties (e.g., choosing El Niño “years” based on season instead of calendar year) allows the model to “learn” about the evolution of the phenomena, thereby reducing the impact of autocorrelation between the training and testing datasets. If data splits do not consider the properties of the phenomenon, then learning relies more on autocorrelation, which by definition inflates skill metrics from the model. Since subseasonal and longer temporal dependencies can result in an insufficient $N_{\mathrm{eff}}$ to represent the processes of interest (e.g., mayer2024), other techniques should be considered. For example, using data from long simulations of climate models to develop the initial versions of the AI/ML models and overcome the limitations of short observational records is becoming more common (e.g., ham2019; riveratello2023). This practice assumes that climate models adequately represent the phenomena of interest and that fine-tuning procedures can correct existing biases. To address spatial autocorrelation, one can stratify geographically distinct regions and build buffer zones between them. If regions are climatologically distinct and such properties are important for model performance, ensuring that data splits contain samples from respective regions while limiting temporal autocorrelation may be appropriate.
@@ 3.1 Working with numerical-valued data: Anomalies and standardization
Most problems in climate prediction involve predicting the anomaly (i.e., departure from an average) in a given field, allowing for skill evaluation beyond just the varying climatology (hamill2006). However, computation of anomalies depends on the problem design, and the computation method can yield different interpretations. A baseline period may be fixed (e.g., 1991–-2020) or centered rolling windows (e.g., $\pm 15$ years), the latter of which may address nonstationarity. Choosing a baseline period early in the record may result in artificial skill gains during evaluation due to the effects of climate change (wulff2022). Choosing a base period later in the record, particularly one with a strong trend, will skew early values. Long-term trends should also be removed to avoid overinflating correlations between input features or relationships derived therein. When calculating anomalies and trends, the spatial element of the climate variables should also be considered. These quantities can be computed using global, regional, or grid cell-specific climatology. A global or regional baseline climatology (latitudinally weighted if using gridded data) may be necessary when spatial anomaly patterns must be preserved (e.g., modes of variability). A grid cell-specific climatology accounts for local variations, which may be useful when interested in strong horizontal gradients (e.g., precipitation anomalies). Altogether, we recommend that there should not be a “one-size-fits-all” approach to anomaly calculation.
@@
A special case exists when using hindcast simulations (i.e., forecasts made with a model for past events). Relatively short hindcast databases (e.g., 10-20 years long) restrict choices for climatologies, potentially leading to averages that use “future” periods and artificially inflating prediction skill (risbey2021). Furthermore, hindcast climatologies are special because they are a function of the initialization date and lead time, so as to avoid model drift and bias. Therefore, when working with hindcast simulations, we suggest computing separate lead time-based climatologies and applying smoothing windows to multiyear averages to address gaps in initialization dates (e.g., pegion2019). For ensemble-based hindcasts, the above-suggested recommendations for anomaly and trend calculations work well for single-model analyses. When considering multi-model means, anomalies should be computed for each model separately.
@@
Finally, feature scaling, either normalization (e.g., min-max scaling) or standardization (i.e., $z$-scoring), of data inputs for AI/ML applications must be carefully done. Single fields with large magnitudes and/or multiple fields with varying magnitudes can cause instability and prevent convergence, requiring feature scaling. We recommend that users consistently scale all features while also considering model design when making such a choice. Feature scaling is unnecessary when using tree-based models (e.g., random forests) since they use feature thresholds in their structure but is especially important when using algorithms sensitive to variance, data containing outliers, and when using distance-based models (e.g., $k$-means clustering). Since outliers can disproportionately affect feature scaling, unless one is interested in predicting outlier events, we suggest either removing the outliers or winsorizing the distribution – i.e., outlier values are reassigned to a specified percentile of the data.
@@
When dealing with variables that have non-Gaussian distributions, other preprocessing transformations can help align input features and should be performed before feature scaling. Fig. 1c provides examples of different transformations done on the distribution of monthly-mean total precipitation maxima in the Indian Monsoon region, depending on user needs: (a) the log transformation (Fig. 1c, red line), (b) the Box-Cox (Fig. 1c, solid blue line; box1964) or Yeo-Johnson (Fig. 1c, dashed red line; yeo2000) transformations, and (c) quantile transformations (Fig. 1c, light blue and solid red lines). Feature scaling should then be applied after these transformations.